# Reuse Your Rewards: Reward Model Transfer for Zero-Shot Cross-Lingual Alignment

### v1

@@ Abstract
[…]
@@ 1 Introduction
We propose a novel reward model (RM) transfer setup, where we re-purpose a trained RM for some source language to align a LM in a target language (Figure 1). Across two tasks (summarization and open-ended dialog generation), two reward optimization methods (reinforcement learning and best-of-$n$ reranking), and various evaluation settings, we demonstrate substantial and consistent zero-shot cross-lingual utility of RMs. Surprisingly, alignment using a different-language RM sometimes outperforms using a same-language RM, both when judged by humans and LMs. We also demonstrate that our RM transfer framework is still useful even when target-language data for supervised finetuning (SFT), another component in alignment, is inaccessible.
@@ 4 Experimental Setup
We consider two tasks: summarization, common in alignment research (Stiennon et al., 2020; Ziegler et al., 2020; Lee et al., 2023; i.a.), and open-ended dialog generation, with substantial real-world relevance. §A describes dataset details and §B training details. §G.1 contains our task instructions.
@@
We use the OpenAssistant dataset (Köpf et al., 2023) with multilingual, pairwise human-rated chat transcripts. For the SFT data, we use the human-preferred response in each pair to finetune the model. Many languages in OpenAssistant have only limited data, so we only consider three languages with the most amounts of data: English, Spanish, and Russian. We use PaLM-2-XXS as the base model (Anil et al., 2023). The authors of OpenAssistant found RL to be ineffective for this dataset (Köpf et al., 2023), which we confirmed in our experiments (Figure 4). We therefore focus on best-of-$n$ for this task.
@@
We assess model quality across several settings. First, we use the target-language RM, which is by design finetuned to judge target-language generation quality. But because of potential RM biases (Gao et al., 2023; Coste et al., 2023; Eisenstein et al., 2023), we also include two zero-shot-prompted evaluation models with much larger backbones—GPT-4 (OpenAI, 2023) and PaLM-2-L (Anil et al., 2023). This latter evaluation setup is common in prior work and has been demonstrated to correlate well with human judgments (Lee et al., 2023; Rafailov et al., 2023; An et al., 2023; Mu et al., 2023; i.a.). We also confirm its validity in §5.1 and §C. Finally, we also perform human evaluations by self-reported native or advanced speakers, though only for a subset of language pairs and 250 (RL) / 100 (best-of-$n$) instances per pair due to its cost. For both human and LM evaluation, we elicit pairwise judgments to compare responses from the aligned model and the SFT model (Bai et al., 2022b; Lee et al., 2023; i.a.). We measure the win rate, i.e., how often the judge prefers the former. A 50% win rate indicates no improvement from alignment. §G.2 includes more details such as the evaluation prompts and positional bias control.
@@ 5.1 Cross-Lingual Alignment Is Effective
Human evaluation (Figure 2) reveals the same trend, though with larger confidence intervals due to the cost. It also justifies LM-based evaluation. For summarization, PaLM-2-L (GPT-4) agrees with humans 65% (69%) of the time in English and 66% (62%) in Spanish, matching the 63% human-human agreement for English reference summaries and 67% for Spanish in Seahorse (Clark, personal communication, April 15, 2024). For dialog, PaLM-2-L (GPT-4) agrees with humans 69% (59%) of the time in English and 62% (60%) in Spanish, again similar to the 63% human-human agreement in Bai et al. (2022a) and 66% in Dubois et al. (2024). With further evidence in §C, we believe our LM judges reasonably reflect output quality.
@@ 5.2 Cross-Lingual Alignment Sometimes Outperforms Monolingual Alignment
This hypothesis is consistent with our observed patterns. First, there are many fewer cases of cross-lingual reward optimization outperforming the monolingual setting when measured by the finetuned target-language RM evaluator than the prompted LM evaluators (Figure 3): under this hypothesis, the finetuned evaluator RMs would be more susceptible to such artifacts and (incorrectly) assign higher scores in the monolingual settings. The underperformance of the translate-train baseline (§5.1) also provides weak evidence: in principle, a source-language RM and a source-translated-into-target-language RM should capture the same reward signal, as they are derived from the same data source, and would lead to similar downstream performance. However, the former is less susceptible to reward over-optimization due to the language mismatch, leading to better performance, though this is admittedly confounded by translation quality.
@@ 5.3 Cross-Lingual Alignment Without Target-Language SFT Data
To account for this, we employ round-trip back-translation, first translating the target-language SFT data into the source language and then back to the target language. This setup is not practically useful but it upper-bounds the effect of translation errors alone. In Figure 5(a), we see that this bridges most of the gap, in some cases leading to models that win over the SFT model >50% of the time. Alternatively, we artificially control for domain by repeating our experiments solely using WikiLingua for both SFT and RM as it is present for all languages. From Figure 5(c), the gap indeed reduces, with the translated SFT models sometimes even outperforming the original, and back-translation is no longer consistently beneficial.
@@ 6.1 Impact of RM Generalizability Within Reward Modeling
Furthermore, the results here do not match their downstream utility, regardless of whether we consider the quality of the RMs as measured by their in-language validation accuracy (Turkish, for example, is the best in Figure 7, but not so in Figure 6), the generalizability of the RMs which we operationalize as the difference between in-language training and validation loss (or accuracy—they yield the same ranking: Russian, German, English, Turkish, Vietnamese, and Spanish, from the least amount of overfitting to the most, again different from Figure 6), or the specific pairwise transfer effectiveness (for each target language, we compare the effectiveness of source languages ranked by the reward modeling task generalizability here vs. by downstream alignment win rate; on summarization, averaged across target languages, Kendall’s $\tau$=0.1 (same when using best-of-$n$ or RL), indicating low ranking agreement). Overall, while cross-lingual alignment depends on RM generalizability on the original task, there are other factors at play too.
@@ Acknowledgments
We would like to thank Jonathan Berant, Jilin Chen, Elizabeth Clark, Daphne Domansi, Jie Fan, Han Guo, Henry Hand, Harrison Lee, Jong Lee, Alisa Liu, Ana Marasović, Usha Rani Markuk, Kathy Meier-Hellstern, Chirag Nagpal, Flavien Prost, Linlu Qiu, Kevin Robinson, Alexis Ross, Shannon Zejiang Shen, Bailin Wang, Xinyan Velocity Yu, and the T5X team Google for their valuable feedback and support. The MIT researchers were partially supported by funds from an MIT-IBM Watson AI Lab grant.
@@ Appendix B Training Details
We use PPO for RL training with a constant learning rate at $10^{-4}$, batch size 32, for 3,000 steps for summarization and 2,500 steps for dialog generation. The value model has 1,000 linear warm-up steps and we only start training the policy model after 2,000 steps elapse. We set the regularization coefficient at $\beta=0.01$.

### v2

@@ Abstract
Zhaofeng ${}^{\text{\Cancer}}$   Ananth ${}^{\text{\Virgo}}$   Yoon ${}^{\text{\Cancer}}$   Jacob ${}^{\text{\Virgo}}$   Ahmad ${}^{\text{\Virgo}}$$^{\text{\Cancer}}$  ${}^{\text{\Virgo}}$oogle DeepMind zfw@csail.mit.edu
@@ 1 Introduction
We propose a novel reward model (RM) transfer setup, where we re-purpose a trained RM for some source language to align a LM in a target language (Figure 1), and investigate the effectiveness of this simple recipe. Across two tasks (summarization and open-ended dialog generation), two reward optimization methods (reinforcement learning and best-of-$n$ reranking), and various evaluation settings, we demonstrate substantial and consistent zero-shot cross-lingual utility of RMs. Surprisingly, alignment using a different-language RM sometimes outperforms using a same-language RM, both when judged by humans and LMs. We also show that our RM transfer framework is useful even when target-language data for supervised finetuning (SFT), another component in alignment, is inaccessible.
@@ 4 Experimental Setup
We consider two tasks: summarization, common in alignment research (Stiennon et al., 2020; Ziegler et al., 2020; Lee et al., 2023; i.a.), and open-ended dialog generation, with substantial real-world relevance. §A describes dataset details and statistics. §B includes training details. §G.1 contains our task instructions.
@@
We use the OpenAssistant dataset (Köpf et al., 2023) with multilingual, pairwise human-rated chat transcripts. [^2] In https://huggingface.co/datasets/OpenAssistant/oasst1. For the SFT data, we use the human-preferred response in each pair to finetune the model. Many languages in OpenAssistant have only limited data, so we only consider three languages with the most amounts of data: English, Spanish, and Russian. We use PaLM-2-XXS as the base model (Anil et al., 2023). The authors of OpenAssistant found RL to be ineffective for this dataset (Köpf et al., 2023), which we confirmed in our experiments (Figure 4). We therefore focus on best-of-$n$ for this task.
@@
We assess model quality across several settings. First, we use the target-language RM, which is by design finetuned to judge target-language generation quality. But because of potential RM biases (Gao et al., 2023; Coste et al., 2023; Eisenstein et al., 2023), we also include two zero-shot-prompted evaluation models with much larger backbones—GPT-4 (OpenAI, 2023) and PaLM-2-L (Anil et al., 2023). This latter evaluation setup is common in prior work and has been demonstrated to correlate well with human judgments (Lee et al., 2023; Rafailov et al., 2023; An et al., 2023; Mu et al., 2023; i.a.). We also confirm its validity in §5.1 and §C. Importantly, both evaluation LMs support multilingual texts. Finally, we also perform human evaluations by self-reported native or advanced speakers, though only for a subset of language pairs and 250 (RL) / 100 (best-of-$n$) instances per pair due to its cost. For both human and LM evaluation, we elicit pairwise judgments to compare responses from the aligned model and the SFT model (Bai et al., 2022b; Lee et al., 2023; i.a.). We measure the win rate, i.e., how often the judge prefers the former. A 50% win rate indicates no improvement from alignment. §G.2 includes more details such as the evaluation prompts and positional bias control.
@@ 5.1 Cross-Lingual Alignment Is Effective
Human evaluation (Figure 2) reveals the same trend, though with larger confidence intervals due to the cost. Human evaluation results also validate and justify LM-based evaluation: For summarization, PaLM-2-L (GPT-4) agrees with humans 65% (69%) of the time in English and 66% (62%) in Spanish, matching the 63% human-human agreement for English reference summaries and 67% for Spanish in Seahorse (Clark, personal communication, April 15, 2024). For dialog, PaLM-2-L (GPT-4) agrees with humans 69% (59%) of the time in English and 62% (60%) in Spanish, again similar to the 63% human-human agreement in Bai et al. (2022a) and 66% in Dubois et al. (2024). With further evidence in §C, we believe our LM judges reasonably reflect output quality.
@@ 5.2 Cross-Lingual Alignment Sometimes Outperforms Monolingual Alignment
This hypothesis is consistent with our observed patterns. First, there are many fewer cases of cross-lingual reward optimization outperforming the monolingual setting when measured by the finetuned target-language RM evaluator than the prompted LM evaluators (Figure 3): under this hypothesis, the finetuned evaluator RMs would be more susceptible to such artifacts and (incorrectly) assign higher scores in the monolingual settings. The underperformance of the translate-train baseline (§5.1) also provides weak evidence: in principle, a source-language RM and a source-translated-into-target-language RM should capture the same reward signal, as they are derived from the same data source, and would lead to similar downstream performance. However, the former is less susceptible to reward over-optimization due to the language mismatch, leading to better performance, though this is confounded by translation quality.
@@ 5.3 Cross-Lingual Alignment Without Target-Language SFT Data
To account for this, we employ round-trip back-translation, first translating the target-language SFT data into the source language and then back to the target language. This setup is not practically useful but it upper-bounds the effect of translation errors alone. Figure 5(a) shows that this bridges most of the gap, sometimes leading to models that win over the SFT model >50% of the time. Alternatively, we control for domain by repeating our experiments solely using WikiLingua for both SFT and RM as it is present for all languages. From Figure 5(c), the gap indeed reduces, with the translated SFT models sometimes even outperforming the original, and back-translation is no longer consistently beneficial.
@@ 6.1 Impact of RM Generalizability Within Reward Modeling
Furthermore, the results here do not match their downstream utility, regardless of whether we consider the quality of the RMs as measured by their in-language validation accuracy (Turkish, for example, is the best in Figure 7, but not so in Figure 6), the generalizability of the RMs which we operationalize as the difference between in-language training and validation loss (or accuracy—they yield the same ranking: Russian, German, English, Turkish, Vietnamese, and Spanish, from the least amount of overfitting to the most, again different from Figure 6), or the specific pairwise transfer effectiveness (for each target language, we compare the effectiveness of source languages ranked by the reward modeling task generalizability here vs. by downstream alignment win rate; on summarization, averaged across target languages, Kendall’s $\tau=0.1$ (same with best-of-$n$ or RL), indicating low ranking agreement). Overall, while cross-lingual alignment depends on RM generalizability on the original task, other factors are at play too.
@@ Acknowledgments
We would like to thank Jonathan Berant, Jilin Chen, Elizabeth Clark, Daphne Domansi, Jie Fan, Han Guo, Henry Hand, Harrison Lee, Jong Lee, Alisa Liu, Ana Marasović, Usha Rani Markuk, Joshua Maynez, Kathy Meier-Hellstern, Chirag Nagpal, Flavien Prost, Linlu Qiu, Kevin Robinson, Alexis Ross, Shannon Zejiang Shen, Bailin Wang, Xinyan Velocity Yu, and the T5X team Google for their valuable feedback and support. The MIT researchers were partially supported by funds from an MIT-IBM Watson AI Lab grant.
@@ Appendix B Training Details
We use PPO for RL training with a constant learning rate at $10^{-4}$, batch size 32, for 3,000 steps for summarization and 2,500 steps for dialog generation. The value model has 1,000 linear warm-up steps and we only start training the policy model after 2,000 steps elapse. We set the regularization coefficient at $\beta=0.1$.